OpenStudy (anonymous):

On a journey a bus takes 35 minutes to travel the first 10 kilometres. It then travels a further 20 kilometres in the next 40 minutes. (a) The bus started the journey at 18 50. At what time did it complete the journey? Answer (a).………………………………… (b) Calculate the average speed of the whole journey in (i) kilometres/minute, Answer (b)(i).……………….………km/min (ii) kilometres/hour. Answer (b)(ii).……………….……..…km/h

OpenStudy (anonymous):

journey completed 20:05

OpenStudy (anonymous):

30/75km/min=2/5km/min=0.4km/min

OpenStudy (anonymous):

c)24km/h